OpenStudy (anonymous):

2x-3y+z=7 x+y+z=4 x+3=2x solve for x, y, and z I got x=23/7 y=73/7 z=22/7 not sure if it is correct

OpenStudy (anonymous):

x+3=2x 3 = 2x - x 3 = x x= 3

OpenStudy (anonymous):

x+y+z=4 3+y+z = 4 y+z = 4-3 y+z = 1 y = 1-z then 2x-3y+z=7 2(3) - 3(1-z) + z = 7 6 -3+3z+z = 7 3+4z = 7 4z = 7-3 4z = 4 z = 4/4 z =1

OpenStudy (anonymous):

So y = 1-z y=1-1 y = 0

OpenStudy (anonymous):

it must be x = 3, y=0, z=1

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you are welcome @typhlosion

OpenStudy (anonymous):

you are the best!

OpenStudy (anonymous):

don't forget for medal :)